In the garden pea the yellow cotyledon is dominant over the green and the thick pod is dominant over the thin pod. When both characters were considered together in self-fertilized dihybrids, the following proportion appeared in the progeny: 100 green, thick pod, 530 yellow, thin pod, 195 yellow, thick pod, 65 green, thin pod. Does this data behave in a 9: 3: 3: 1 ratio?
Answer #1

Answer:

Answer:

Yellow = Y ; Green = y

Thick pod = P ; Thin pod = p

YyPp x YyPp ----Parents (Self fertilization of Dihybrid)

Yellow cotyledon, thick pod = 9/16

Yellow cotyledon, thin pod = 3/16

Green cotyledon, thick pod = 3/16

Green cotyledon, thin pod = 1/16

Phenotype Observed(O) Expected (E) O-E (O-E)2 (O-E)2/E
yellow, thick 195 500.63 -305.625 93406.64 186.58
Yellow, thin 530 166.875 363.125 131859.77 790.17
Green, thick 100 166.875 -66.875 4472.27 26.80
Green, thin 65 55.625 9.375 87.89 1.58
Total 890 890 1005.13

Chi-square value = 1005.13

Degrees of freedom = number of categories - 1

= 4-1 = 3

The chi-square value of 1005.13 is greater than the critical value of 7.81 (3 degrees of freedom at 5% p value). Hence we reject null hypothesis and the data is not fit in 9:3:3:1 ratio